The heigth of person is 400cm. Express it into m and mm

Grade:6

2 Answers

Arun
4 years ago
Dearr student
 
As we know 1 cm = 0.01 m
And 1 cm = 10 mm
 
Hence 400 cm = 400 * 0.01 = 4 m
And 400cm = 400 * 10 = 4000 mm
KALYAN
3 years ago
to convert centmeters to meters and centimeters to millimeters we need to we need to how many centimeters is equal to meters and how many millimeters is equal to centimeter
 
 
100 cm = 1 m
10 mm = 1 cm
 
so,
400cm =400/100 =4 m
400cm =400*10 =4000 mm
 
approve my answer if you liked it.